# PolyNet: Learning Diverse Solution Strategies for Neural Combinatorial Optimization

### v1

@@ Abstract
Reinforcement learning-based methods for constructing solutions to combinatorial optimization problems are rapidly approaching the performance of human-designed algorithms. To further narrow the gap, learning-based approaches must efficiently explore the solution space during the search process. Recent approaches artificially increase exploration by enforcing diverse solution generation through handcrafted rules, however, these rules can impair solution quality and are difficult to design for more complex problems. In this paper, we introduce PolyNet, an approach for improving exploration of the solution space by learning complementary solution strategies. In contrast to other works, PolyNet uses only a single-decoder and a training schema that does not enforce diverse solution generation through handcrafted rules. We evaluate PolyNet on four combinatorial optimization problems and observe that the implicit diversity mechanism allows PolyNet to find better solutions than approaches the explicitly enforce diverse solution generation.
@@ 1 Introduction
A key limitation of learning-based approaches is that they often struggle to sufficiently explore the solution space, which is known to be a critical aspect of metaheuristic optimization algorithms (Blum & Roli, 2003). Many recent neural construction methods (e.g., Li et al. (2023); Choo et al. (2022)) follow the POMO approach (Kwon et al., 2020) and improve exploration by forcing diverse first actions during solution construction. This assumes that the first construction action has a negligible impact on solution quality, which is true for the traveling salesperson problem (TSP) and the capacitated vehicle routing problem (CVRP). However, in more complex optimization problems the initial action often significantly influences solution quality, rendering these methods less effective at generating solutions. Hence, many recent approach designed only for the TSP and CVRP do not seamlessly extend to other optimization problems.
Additionally to (or instead of) increasing exploration by forcing the initial construction action, diversity has been increased by using multi-decoder models. Xin et al. (2021) propose a transformer model with multiple decoders that encourage each decoder to learn a distinct solution strategy during training by maximizing the Kullback-Leibler (KL) divergence between decoder output probabilities. However, to manage computational costs, diversity is only promoted in the initial construction step. In contrast, Grinsztajn et al. (2023) introduce Poppy, a training procedure for multi-decoder models that increases diversity without relying on KL divergence. Poppy generates a population (i.e., a set) of decoders during the learning phase by training only the best-performing decoder for each problem instance. While effective, Poppy is computationally intensive, requiring a separate decoder for each policy, thus limiting the number of learnable policies per problem.
@@
By utilizing a single decoder to learn multiple strategies, PolyNet quickly generates a set of diverse solutions for a problem instance. This significantly enhances exploration, allowing us to find better solution during training and testing. Furthermore, by abandoning the concept of forcing diverse first actions, we exclusively rely on PolyNet’s inherent diversity mechanism to facilitate exploration during the search process. This fundamental change significantly impacts solution generation for problems in which the first move greatly influences performance, such as in the CVRP with time windows (CVRPTW).
We assess PolyNet’s performance across four problems: the TSP, the CVRP, CVRPTW, and the flexible flow shop problem (FFSP). Across all problems, PolyNet consistently demonstrates a significant advancement over the state-of-the-art in both swift solution generation and comprehensive search efforts. Moreover, we observe that the solution diversity arising during PolyNet’s training enables the discovery of superior solutions compared to artificially enforcing diversity by fixing the initial solution construction step.
@@ 2 Literature Review
Neural CO In their seminal work, Vinyals et al. (2015) introduce the pointer network architecture, an early application of modern machine learning methods to solve CO problems. Pointer networks autoregressively generate discrete outputs corresponding to input positions. When trained via supervised learning, they can produce solutions for the TSP with up to $50$ nodes. Bello et al. (2016) propose training pointer networks using reinforcement learning instead and illustrate the efficacy of this method in solving larger TSP instances.
@@
Nazari et al. (2018) use the pointer network architecture to solve the CVRP with 100 nodes. Kool et al. (2019) improve this architecture by using a transformer-based encoder with self-attention (Vaswani et al., 2017). Recognizing that many CO problems contain symmetries, Kwon et al. (2020) propose POMO, a method that leverages symmetries to generate better solutions. Kim et al. (2022) extend these ideas and propose a general-purpose symmetric learning scheme. Drakulic et al. (2023) use bisimulation quotienting (Givan et al., 2003) to improve out-of-distribution generalization of neural CO methods. Few works, e.g. Falkner & Schmidt-Thieme (2020) and Kool et al. (2022a), propose neural CO approaches for routing problems with time windows.
Instead of constructing solutions autoregressively, some methods predict a heat-map that emphasize promising solution components (e.g., arcs in a graph). This is subsequently used in a post-hoc search for solution construction (Joshi et al., 2019; Fu et al., 2021; Kool et al., 2022b). Another class of methods iteratively improves initial yet complete solutions. For instance, Hottung & Tierney (2020) propose a framework that iteratively destroys parts of a solution using handcrafted procedures and then repairs it using a learned operator to explore the solution space. Similarly, Ma et al. (2021) learn to iteratively improve an initial solution by performing local adjustments, while Chen & Tian (2019) pose CO as a sequential rewriting problem in which parts of a solution are iteratively changed using a learned routine.
@@
Hottung et al. (2022) introduce efficient active search (EAS), a method that guides the search by updating a subset of the policy parameters during inference. Choo et al. (2022) propose SGBS, a post-hoc inference mechanism that combines Monte-Carlo tree search with beam search to provide search guidance. When used in combination with EAS, it achieves state-of-the-art performance on several problems.
Diversity Mechanisms in RL In RL, skill-learning algorithms aim to discover a set of policies with diverse behaviors (defined by the visited states) to accelerate task-specific training. For instance, Eysenbach et al. (2019) and Sharma et al. (2019) learn skills that exhibit predictable behavior and are as diverse as possible. Here, a skill corresponds to the policy conditioned on some latent context.
In contrast to implicitly maintaining a collection of agents through context, population-based RL techniques explicitly maintain a finite agent population and use diversity mechanisms to discover diverse strategies for solving RL tasks. For example, Zhang et al. (2019) learn a collection of policies that solve a task using distinct action sequences. Alternatively, agents can be trained in a population-based setup (Pierrot & Flajolet, 2023). Recently, Wu et al. (2023) focus on task-specific diversity defined according to user-specified behavior descriptors and employ population-based training to maximize diversity at different quality levels.
@@
Diversity mechanisms are also used in single and multi-agent setups to learn diverse problem-solving strategies. For instance, in neural program synthesis, Bunel et al. (2018) optimize the expected reward when sampling a pool of solutions and keep the best one. This encourages the policy to diversify its choices by assigning probability mass to several solutions. In multi-agent RL, Li et al. (2021) use the mutual information between agents’ identities and trajectories as an intrinsic reward to promote diversity and thus solve cooperative tasks requiring diverse strategies.
@@
Diversity Mechanisms in Neural CO Kim et al. (2021) present a hierarchical strategy for solving routing problems, where a learned “seeder policy” maximizes solution diversity through entropy rewards and generates diverse candidate solutions. These are then modified by a learned “reviser policy” to improve solution quality. As previously discussed, Xin et al. (2021) encourage diverse solutions using multiple decoders and KL divergence regularization, while Grinsztajn et al. (2023) use a population of agents through multiple decoders to learn complementary strategies, updating exclusively the best-performing agent at each iteration. This objective shares similarities with Bunel et al. (2018), in which also only the best solution is considered. Hottung et al. (2021) and Chalumeau et al. (2023) learn a continuous latent space that encodes solutions for CO problems and that can be used to sample diverse solutions at test time.
@@ 3.1 Background
Neural CO approaches seek to train a neural network denoted as $\pi_{\theta}$ with learnable weights $\theta$. The network’s purpose is to generate a solution $\tau$ when provided with an instance $l$. To achieve this, we employ RL techniques and model the problem as a Markov decision process (MDP), wherein a solution is sequentially constructed in $T$ discrete time steps. At each step $t\in(1,\dots,T)$, an action $a_{t}$ is selected based on the probability distribution $\pi_{\theta}(a_{t}|s_{t})$ defined by the neural network where $s_{t}$ is the current state. The initial state $s_{1}$ encapsulates the information about the problem instance $l$, while subsequent states $s_{t+1}$ are derived by applying the action $a_{t}$ to the previous state $s_{t}$. A (partial) solution denoted as $\bar{\tau}_{t}$ is defined by the sequence of selected actions $(a_{1},a_{2},\dots,a_{t})$. Once a complete solution $\tau=\bar{\tau}_{T}$ satisfying all problem constraints is constructed, we can compute its associated reward ${\mathcal{R}}(\tau,l)$. The overall policy of generating a solution $\tau$ for an instance $l$ is defined as $\pi_{\theta}(\tau\mid l)=\prod_{t=1}^{T}\pi_{\theta}(a_{t}\mid s_{t})$.
@@ 3.2 Overview
PolyNet is a learning-based approach designed to learn a set of diverse solution strategies for CO problems. During training, each strategy is allowed to specialize on a subset of the training data, and thus need not be the best strategy for the entire dataset. This essentially results in a portfolio of strategies, which are known to be highly effective for solving CO problems (Bischl et al., 2016). Our pursuit of diversity is fundamentally a means to enhance exploration and consequently solution quality. Note that PolyNet not only enhances performance at test time (where we sample multiple solutions for each strategy and keep only the best one), but also improves exploration during training.
@@ 3.3 Training
where $b_{\circ}$ is a baseline (which we discuss in detail below). Updating the model weights only based on the best found solution allows the network to learn specialized strategies that do not have to work well on all instances from the training set. While this approach does not explicitly enforce diversity, it incentivizes the model to learn diverse strategies in order to optimize overall performance. For a more in depth discussion on how this loss leads to diverse solution generation we refer the reader to Grinsztajn et al. (2023).
@@
Many recent neural construction heuristics follow the POMO approach and rollout $N$ solutions from $N$ different starting states per instance to increase exploration. This is possible because many CO problems contain symmetries in the solution space that allow an optimal solution to be found from different states. In practice, this mechanism is implemented by forcing a different first construction action for each of the $N$ rollouts. Forcing diverse rollouts not only improves exploration, but also allows the average reward of all $N$ rollouts to be used as a baseline. However, this exploration mechanism should not be used when the first action can not be freely chosen without impacting the solution quality.
In PolyNet, we do not use an exploration mechanism or a baseline that assumes symmetries in the solution space during training. Instead, we only rely on the exploration provided by our conditional solution generation. This allows us to solve a wider range of optimization problems. As a baseline we simply use the average reward of all $K$ rollouts for an instance, i.e., $b_{\circ}=\frac{1}{K}\sum_{i=1}^{K}R(\tau_{i},l)$.
@@ 3.5 Search
To facilitate a more extensive, guided search, PolyNet can be combined with EAS. EAS is a simple technique to fine-tune a subset of model parameters for a single instance in an iterative process driven by gradient descent. In contrast to the EAS variants described in Hottung et al. (2022), we do not insert any new layers into the network or update the instance embeddings. Instead, we only fine-tune the new PolyNet layers during search. Since PolyNet is specifically trained to create diverse solutions based on these layers, EAS can easily explore a wide variety of solutions while modifying only a portion of the model’s parameters.
@@ 4 Experiments
We compare PolyNet’s search performance to state-of-the-art methods on four established problems. We also explore solution diversity during the training and testing phases and analyze the impact of the free first move selection. An ablation study for our network architecture changes is provided in Appendix B. While most of our evaluation is focused on routing problems, we demonstrate that PolyNet can also be used to solve other CO problems by applying PolyNet to the FFSP (see Appendix A). Our experiments are conducted on a GPU cluster utilizing a single Nvidia A100 GPU per run. In all experiments, the new PolyNet layers comprise two linear layers, each with a dimensionality of $256$. PolyNet will be made publicly available upon acceptance.

### v2

@@ Abstract
Reinforcement learning-based methods for constructing solutions to combinatorial optimization problems are rapidly approaching the performance of human-designed algorithms. To further narrow the gap, learning-based approaches must efficiently explore the solution space during the search process. Recent approaches artificially increase exploration by enforcing diverse solution generation through handcrafted rules, however, these rules can impair solution quality and are difficult to design for more complex problems. In this paper, we introduce PolyNet, an approach for improving exploration of the solution space by learning complementary solution strategies. In contrast to other works, PolyNet uses only a single-decoder and a training schema that does not enforce diverse solution generation through handcrafted rules. We evaluate PolyNet on four combinatorial optimization problems and observe that the implicit diversity mechanism allows PolyNet to find better solutions than approaches that explicitly enforce diverse solution generation.
@@ 1 Introduction
A key limitation of learning-based approaches is that they struggle to sufficiently explore the solution space. The metaheuristics literature identifies exploration as a key component of heuristic optimization procedures (Blum & Roli, 2003), but examining diverse solutions alone is not enough to find high-quality solutions. Mahfoud (1995) discuss the need for useful diversity, i.e., diversity that “helps in achieving some purpose or goal.” To encourage search diversity, many recent neural construction methods (e.g., Li et al. (2023a); Choo et al. (2022)) follow the POMO approach (Kwon et al., 2020) and improve exploration by forcing diverse first actions during solution construction. While this creates significant diversity through exploiting symmetries in the problem representation (every solution has a different first action), no effort is made to encourage diversity subsequent to the first action. Furthermore, in more complex optimization problems the initial action can significantly influence the solution quality, rendering these methods less effective at generating solutions.
Acknowledging that useful diversity is an essential component of search techniques, neural CO approaches have begun attempting to encourage more exploration during search. Xin et al. (2021) propose a transformer model with multiple decoders that encourage each decoder to learn a distinct solution strategy during training by maximizing the Kullback-Leibler (KL) divergence between decoder output probabilities. However, to manage computational costs, diversity is only promoted in the initial construction step. In contrast, Grinsztajn et al. (2023) introduce Poppy, a training procedure for multi-decoder models that increases diversity without relying on KL divergence. Poppy generates a population (i.e., a set) of decoders during the learning phase by training only the best-performing decoder for each problem instance. While effective, Poppy is computationally intensive, requiring a separate decoder for each policy, thus limiting the number of learnable policies per problem.
@@
By utilizing a single decoder to learn multiple strategies, PolyNet quickly generates a set of diverse solutions for a problem instance. This significantly enhances exploration, allowing us to find better solution during training and testing. Furthermore, by abandoning the concept of forcing diverse first actions, we exclusively rely on PolyNet’s inherent diversity mechanism to facilitate exploration during the search. This fundamental change improves solution generation for problems in which the first move greatly influences performance, such as in the CVRP with time windows (CVRPTW).
PolyNet is an autoregressive construction method, which makes a direct application to large-scale routing problems computationally impractical. However, small- and medium-scale routing problems, which often involve numerous constraints (e.g., customer time windows), are frequently encountered in practice. This is reflected in the current focus of the operations research community, which prioritizes smaller problems with numerous complex constraints (e.g., Zhang et al. (2024); Cataldo-Díaz et al. (2024); Lera-Romero et al. (2024); Jolfaei & Alinaghian (2024)). With PolyNet, we aim to take a step toward addressing these more complex problems by enabling diverse solution generation without relying on the symmetries characteristic of simpler routing problems.
We assess PolyNet’s performance across four problems: the TSP, the CVRP, CVRPTW, and the flexible flow shop problem (FFSP). Across all problems, PolyNet consistently demonstrates a significant advancement over the state-of-the-art. Moreover, we observe that the solution diversity arising during PolyNet’s training enables the discovery of superior solutions compared to artificially enforcing diversity by fixing the initial solution construction step.
@@ 2 Literature review
Neural CO In their seminal work, Vinyals et al. (2015) introduce the pointer network architecture, an early application of modern machine learning methods to solve CO problems. Pointer networks autoregressively generate discrete outputs corresponding to input positions. When trained via supervised learning, they can produce solutions for the TSP with up to $50$ nodes. Bello et al. (2016) propose training pointer networks using reinforcement learning instead and illustrate the efficacy of this method in solving larger TSP instances. However, these early methods are limited in generalization and in the quality of the generated solutions.
@@
Nazari et al. (2018) extend the pointer network architecture to address the CVRP with up to 100 nodes. Building on this, Kool et al. (2019) enhance the architecture by introducing a transformer-based encoder with self-attention (Vaswani et al., 2017), enabling the model to learn more effective policies and produce higher-quality solutions. Recognizing that many CO problems exhibit inherent symmetries, Kwon et al. (2020) propose POMO, a method that exploits these symmetries to generate diverse solutions. However, POMO’s effectiveness is heavily dependent on the assumption that such symmetries can be efficiently leveraged, which limits its applicability to simple problems. Kim et al. (2022) extend these ideas and propose a general-purpose symmetric learning scheme. Drakulic et al. (2023) use bisimulation quotienting (Givan et al., 2003) to improve out-of-distribution generalization of neural CO methods. However, they use imitation learning which requires a large training set of high-quality solutions limiting the applicability of their method. Luo et al. (2023) improves generalization in neural CO by using a “heavy” decoder to make more effective node selections during solution construction, but this increases computational overhead. Despite these advancements, only a few works, such as Falkner & Schmidt-Thieme (2020), Kool et al. (2022a), Hua et al. (2025), and Berto et al. (2024b), target routing problems with additional constraints like time windows, highlighting a gap in extending neural CO approaches to more complex problem variants.
Instead of constructing solutions autoregressively, some methods predict a heat-map that emphasizes promising solution components (e.g., arcs in a graph) and is subsequently used in a post-hoc search for solution construction (Joshi et al., 2019; Fu et al., 2021; Kool et al., 2022b; Min et al., 2023). However, Xia et al. (2024) reveal that these methods often underperform compared to simpler baselines. Another class of methods iteratively improves initial solutions. Hottung & Tierney (2020), Ma et al. (2021), Chen & Tian (2019), and Hottung et al. (2025) iteratively improve an initial solution by performing local adjustments. Similarly, several works guide the $k$-opt heuristic for vehicle routing problems (Wu et al., 2019; da Costa et al., 2020; Ma et al., 2023). These improvement methods require longer runtimes to be effective. In contrast, PolyNet also excels at generating high-quality solutions quickly. Sun & Yang (2023) cast CO problems as discrete binary vector optimization problems and utilize denoising diffusion models for solution generation and Kim et al. (2024) utilize the solution symmetry in the combinatorial space to improve the sample efficiency during training. Both approaches fail to generalize to problems with more complex constraints like the CVRPTW.
@@
Hottung et al. (2022) introduce efficient active search (EAS), a method that guides the search by updating a subset of the policy parameters during inference. Choo et al. (2022) propose SGBS, an inference mechanism that combines Monte-Carlo tree search with beam search to provide search guidance. When used in combination with EAS, it achieves state-of-the-art performance on several problems. Li et al. (2023b) learn a distribution of high-quality solutions using diffusion models and perform a gradient-based search at test time. PolyNet outperforms these recent approaches while being able to even tackle problems with more complex constraints.
Diversity in metaheuristics In the context of metaheuristics, solution diversity plays a crucial role in balancing the exploration-exploitation trade-off. Population-based algorithms, such as genetic algorithms, aim to maintain a pool of diverse solutions to promote thorough exploration of the search space (Gendreau et al., 2010). Additionally, some works specifically target generating multiple solutions for single-objective problems. For instance, Hanaka et al. (2022) introduce a polynomial-time algorithm to compute multiple diverse shortest paths in weighted directed graphs. Similarly, Huang et al. (2019) propose a niching memetic algorithm designed to identify multiple optimal solutions for TSP instances, emphasizing diversity within the solution set.
Diversity mechanisms in RL Diversity mechanisms in RL aim to encourage exploration by generating a range of strategies or solutions. Skill-learning algorithms, such as those by Eysenbach et al. (2018) and Sharma et al. (2019), aim to learn policies with diverse behaviors to accelerate training. However, these methods are not directly applicable to CO tasks, where diversity must also consider the quality of solutions, not just behavioral variation.
@@
In contrast to implicitly maintaining a collection of agents through context, population-based RL techniques explicitly maintain a finite agent population and use diversity mechanisms to discover diverse strategies for solving RL tasks, e.g., Zhang et al. (2019), Pierrot & Flajolet (2023), Wu et al. (2023). In neural program synthesis, Bunel et al. (2018) optimize the expected reward when sampling a pool of solutions and keep the best one. Li et al. (2021) use the mutual information between agents’ identities and trajectories as an intrinsic reward to promote diversity and thus solve cooperative tasks requiring diverse strategies.
@@
Diversity mechanisms in neural CO Kim et al. (2021) present a hierarchical strategy for solving routing problems, where a base policy learns to generate diverse candidate solutions through entropy regularization. While entropy regularization can be used to increase diversity during solution sampling, it does not allow to learn truly different strategies. Xin et al. (2021) encourage diverse solutions using multiple decoders and KL divergence regularization, while Grinsztajn et al. (2023) use an agent population through multiple decoders to learn complementary strategies, updating exclusively the best-performing agent at each iteration similar to Bunel et al. (2018). While using multiple-decoders encourages diversity, it results in a large computational overhead. Hottung et al. (2021) and Chalumeau et al. (2023) learn a continuous latent space that encodes solutions for CO problems and use it to sample diverse solutions at test time. Both approaches are designed for longer runtimes and can not be easily applied to more difficult optimization problems.
@@ 3.1 Background
Neural CO approaches seek to train a neural network denoted as $\pi_{\theta}$ with learnable weights $\theta$. The network’s purpose is to generate a solution $\tau$ when provided with an instance $l$. To achieve this, we employ RL techniques and model the problem as a Markov decision process (MDP), wherein a solution is sequentially constructed in $T$ discrete time steps. At each step $t\in(1,\dots,T)$, an action $a_{t}$ is selected based on the probability distribution $\pi_{\theta}(a_{t}|s_{t})$ defined by the neural network where $s_{t}$ is the current state. The initial state $s_{1}$ encapsulates the information about the problem instance $l$, while subsequent states $s_{t+1}$ are derived by applying the action $a_{t}$ to the previous state $s_{t}$. A (partial) solution denoted as $\bar{\tau}_{t}$ is defined by the sequence of selected actions $(a_{1},a_{2},\dots,a_{t})$. Once a complete solution $\tau=\bar{\tau}_{T}$ satisfying all problem constraints is constructed, we can compute its associated reward ${\mathcal{R}}(\tau,l)$. The overall probability of generating a solution $\tau$ for an instance $l$ is defined as $\pi_{\theta}(\tau\mid l)=\prod_{t=1}^{T}\pi_{\theta}(a_{t}\mid s_{t})$.
@@ 3.2 Overview
PolyNet is a learning-based approach designed to learn a set of diverse solution strategies for CO problems. During training, each strategy is allowed to specialize on a subset of the training data. This essentially results in a portfolio of strategies, which are known to be highly effective for solving CO problems (Bischl et al., 2016). Our pursuit of diversity is fundamentally a means to enhance exploration and consequently solution quality. Note that PolyNet not only enhances performance at test time (where we sample multiple solutions for each strategy and keep only the best one), but also improves exploration during training.
@@ 3.3 Training
where $b_{\circ}$ is a baseline (which we discuss in detail below). Updating the model weights only based on the best found solution allows the network to learn specialized strategies that do not have to work well on all instances from the training set. While this approach does not explicitly enforce diversity, it incentivizes the model to learn diverse strategies in order to optimize overall performance. For a more in depth discussion on how this loss leads to diverse solution generation, see Grinsztajn et al. (2023).
@@
Many recent neural construction heuristics follow the POMO approach and rollout $N$ solutions from $N$ different starting states per instance to increase exploration. This is possible because many CO problems contain symmetries in the solution space that allow an optimal solution to be found from different states. In practice, this mechanism is implemented by forcing a different first construction action for each of the $N$ rollouts. Forcing diverse rollouts significantly improves exploration. However, this exploration mechanism should not be used when the first action can not be freely chosen without impacting the solution quality.
In PolyNet, we do not use an exploration mechanism that assumes symmetries in the solution space during training. Instead, we only rely on the exploration provided by our conditional solution generation. This allows us to solve a wider range of optimization problems. As a baseline we simply use the average reward of all $K$ rollouts for an instance, i.e., $b_{\circ}=\frac{1}{K}\sum_{i=1}^{K}R(\tau_{i},l)$.
Figure 2: Decoder architecture.
@@ 3.5 Search
To facilitate a more extensive, guided search, PolyNet can be combined with EAS. EAS is a simple technique to fine-tune a subset of model parameters for a single instance in an iterative process driven by gradient descent. In contrast to the EAS variants described in Hottung et al. (2022), we do not insert any new layers into the network or update the instance embeddings. Instead, we only fine-tune the new PolyNet layers during search. Since PolyNet is specifically trained to create diverse solutions based on these layers, EAS can easily explore a wide variety of solutions while modifying only a subset of the model parameters.
@@ 4 Experiments
We compare PolyNet’s search performance to state-of-the-art methods on four established problems. We also explore solution diversity during the training and testing phases and analyze the impact of the free first move selection. An ablation study for our network architecture changes is provided in Appendix B. While most of our evaluation is focused on routing problems, we demonstrate that PolyNet can also be used to solve other CO problems by applying PolyNet to the FFSP (see Appendix A). Our experiments are conducted on a GPU cluster utilizing a single Nvidia A100 GPU per run. In all experiments, the new PolyNet layers comprise two linear layers, each with a dimensionality of $256$. Our implementation of PolyNet is available at https://github.com/ahottung/PolyNet. Furthermore, PolyNet is also implemented in the RL4CO framework (Berto et al., 2024a).